# Truth or Twist? Optimal Model Selection for Reliable Label Flipping Evaluation in LLM-based Counterfactuals

### v1

@@ Abstract
Counterfactual examples are widely employed to enhance the performance and robustness of large language models (LLMs) through counterfactual data augmentation (CDA). However, the selection of the judge model used to evaluate label flipping, the primary metric for assessing the validity of generated counterfactuals for CDA, yields inconsistent results. To decipher this, we define four types of relationships between the counterfactual generator and judge models. Through extensive experiments involving two state-of-the-art LLM-based methods, three datasets, five generator models, and 15 judge models, complemented by a user study $(n=90)$, we demonstrate that judge models with an independent, non-fine-tuned relationship to the generator model provide the most reliable label flipping evaluations. [^1] Our code will be publicly available once the paper is published. Relationships between the generator and judge models, which are closely aligned with the user study for CDA, result in better model performance and robustness. Nevertheless, we find that the gap between the most effective judge models and the results obtained from the user study remains considerably large. This suggests that a fully automated pipeline for CDA may be inadequate and requires human intervention.
@@ 1 Introduction
Given superior capabilities of LLMs across a wide range of natural language processing tasks, they are convincingly employed to generate counterfactual examples through well-designed prompts Chen et al. (2023); Li et al. (2024). Counterfactual examples are minimally altered versions of original inputs that flip the initial label Miller (2019); Ross et al. (2021); Madsen et al. (2022). They serve as a valuable approach for probing black-box LLMs Wang et al. (2024) and for CDA aimed at improving model robustness and performance [^2] We want to emphasize the subtle yet significant distinction between counterfactuals used for explaining model predictions and those used for CDA. In the former, the objective is to flip the model’s prediction, while the goal of CDA is to flips the ground truth label (Figure 2). In this study, we mainly focus on the latter type of counterfactuals used for CDA. Dixit et al. (2022); Balashankar et al. (2023); Agrawal et al. (2025). To evaluate the effectiveness and validity of the LLM-generated counterfactuals for CDA, the label flip rate (LFR) is a common metric of choice Ge et al. (2021). It is the percentage of valid counterfactuals where the ground truth labels are flipped out of the total number of instances. LFR of counterfactuals can be evaluated using either the same model that generates the counterfactual Bhattacharjee et al. (2024a, b); Wang et al. (2025) or independent models Dixit et al. (2022); Balashankar et al. (2023). The optimal strategy for selecting models to evaluate the ground-truth validity of counterfactuals for CDA remains uncertain, which, in turn, constrains efforts to enhance model robustness and performance through CDA.
@@
In this work, we first define four types of relationships between the counterfactual generator model and the judge model: being the same model, independent models with and without fine-tuning on the target dataset, distilled models, and models from the same family. Secondly, we conduct comprehensive experiments to predict labels for counterfactuals generated by two state-of-the-art approaches across three datasets and five generator models, with 15 judge models. Thirdly, we undertake a user study to assess the validity of the generated counterfactuals in acquiring ground-truth LFR. Our finding shows that a judge model with an independent, non-fine-tuned relationship to the generator captures label flipping most effectively. Relationships between the generator and judge models that are most aligned with the user study, lead to improved model performance and robustness. Additionally, there remains a considerable gap between the performance of the best judge models and the results observed in the user study.
2 Related Work
Counterfactual Evaluation
Next to LFR3), counterfactual examples can be assessed by the following key aspects: Similarity measures the edit distance required to generate the counterfactual Wu et al. (2021); Kaushik et al. (2020); Nguyen et al. (2024b). Diversity quantifies the average pairwise distance between multiple counterfactual examples for a given input Wu et al. (2021); Chen et al. (2023). Fluency describes the similarity of a counterfactual example to human-written text Robeer et al. (2021); Madaan et al. (2021). [^3] Besides, counterfactuals can be evaluated through human judgment or LLM-as-a-Judge based on such as coherence, understandability, feasibility, fairness, and completeness Nguyen et al. (2024a); Domnich et al. (2025). While automated metrics exist for other aspects, evaluating LFR for CDA requires ground-truth labels and human judgment. In this paper, we thoroughly investigate which type of relationship between the generator and judge model is preferable for verifying label flipping based on the results of the user study.
@@ 2 Related Work
Prior to, and even during the widespread adoption of LLMs, encoder-only models, e.g., DeBERTa Dixit et al. (2022), RoBERTa Ross et al. (2021); Balashankar et al. (2023) or BERT Kaushik et al. (2020); Fern and Pope (2021); Robeer et al. (2021), were predominantly used to verify label flipping. This preference stems from their advanced performance in text classification tasks. In more recent work, the same LLMs are increasingly employed both to generate counterfactuals Bhattacharjee et al. (2024a, b); Wang et al. (2025); Dehghanighobadi et al. (2025) and to evaluate them for label flipping, since their classification accuracy rivals that of fine-tuned encoder-only models.
@@ 3.1 Label Flipping
Given that counterfactual examples used for CDA are defined as edited inputs that alter ground truth labels, LFR is positioned as the primary evaluation metric for assessing the effectiveness and validity of generated counterfactuals Ross et al. (2021); Nguyen et al. (2024a); Wang et al. (2025). LFR is quantified as the percentage of instances in which labels are successfully flipped relative to the total number of counterfactuals, where $N$ stands for the total number of counterfactuals, $y_{k}$ represents the ground-truth label of the original input, $y_{k}^{{}^{\prime}}$ denotes the prediction of its corresponding counterfactual and $\mathds{1}$ is the indicator function:
@@ 3.2 Relationships
To determine the optimal model selection strategy for label flip evaluation, we identified four prevalent relationships $\mathcal{R}(\mathcal{G},\mathcal{J})$ between the counterfactual generator model $LLM_{\mathcal{G}}$ and judge models $LLM_{\mathcal{J}}$, which are used to assess label flipping:
  • Same model ($\mathcal{R}_{sm}$): the two models are identical.
@@
  • Independent models ($\mathcal{R}_{im}$): the two models belong to different model families. We further distinguish the independent models based on whether $LLM_{\mathcal{J}}$ is fine-tuned on the given dataset: independent models with ($\mathcal{R}_{imw}$) and without ($\mathcal{R}_{imwo}$) fine-tuning.
@@ 4.1 Counterfactual Methods Selection
We select two state-of-the-art approaches based on LLMs that are shown to generate counterfactual examples efficiently and effectively: FIZLE Bhattacharjee et al. (2024a) and FLARE Bhattacharjee et al. (2024b). FIZLE first enables LLMs to identify key words within the input and then leverages these words to guide the generation of counterfactual examples. Meanwhile, FLARE generates counterfactuals by prompting LLMs in three steps: extracting latent features, identifying relevant words linked to those features, and minimally modifying these words to produce counterfactual examples.
@@ 4.2 Datasets
We use three widely studied classification tasks for counterfactual generation in the literature with varying complexity [^4] Examples of the dataset and the label distribution are included in Appendix A.:
@@ 4.3 Models
We select five instruction-tuned LLMs varying in parameter size and employ them to generate counterfactuals: Qwen2.5-{7B,14B,32B} Qwen (2024), Llama3-{8B,70B} AI@Meta (2024), which also serve as judge models $LLM_{\mathcal{J}}$, considered as $\mathcal{R}_{sm}$ relationship (§3.2). Additionally, we deploy fine-tuned BERT Devlin et al. (2019) and RoBERTa Liu et al. (2020) on the target datasets (§4.2), along with out-of-the-box Phi4-14B Abdin et al. (2024), Qwen2.5-72B Qwen (2024), Mistral-Large-Instruct Jiang et al. (2023), and DeepSeek-R1-Distill-{Qwen,Llama} DeepSeek-AI (2025), as $LLM_{\mathcal{J}}$ [^5] Detailed information about the models employed is provided in Appendix B, and the downstream task performance of each model across three datasets is presented in Appendix C. (Table 1). Since $LLM_{\mathcal{J}}$ are used to identify label flipping, we evaluate their downstream task performance in terms of classification accuracy across three datasets. Table 3 illustrates that $LLM_{\mathcal{J}}$ with $\mathcal{R}_{imw}$ relationship, generally achieve the highest downstream performance.
@@ 4.4 Automatic Evaluation
LFR is evaluated based on the classification results of counterfactual examples generated using FIZLE Bhattacharjee et al. (2024a) and FLARE Bhattacharjee et al. (2024b) [^6] The classification prompts used are shown in Figure 6., using the deployed judge models $LLM_{\mathcal{J}}$ described in Section 4.3.
4.5 User Study
We conduct a user study with 90 native English-speaking participants. We randomly select 45 dataset indices for each of the three datasets. For each subset, i.e., generator-dataset pair (Table 1), the counterfactuals generated by the corresponding generator model $LLM_{\mathcal{G}}$ for the selected indices are evaluated by at least two human annotators. [^7] If no majority label emerges from the labels provided by human annotators, we manually determine which label should be selected. The annotation guidelines and annotator information are provided in Appendix D. We further conduct an automatic evaluation in Appendix F on the selected 45 counterfactuals as a sanity check to validate their representativeness of the overall distribution. Each annotator will be given 15 counterfactuals, along with the set of possible labels given by the dataset (§4.2), and tasked with selecting the optimal label. We report an inter-annotator agreement of Cohen’s $\kappa=0.55$. Lastly, we calculate the ground-truth LFR as the proportion of valid counterfactuals relative to the total number of instances.
Model Judge Model ($LLM_{\mathcal{J}}$) AG News SST2 SNLI
Qwen2.5-7B Qwen2.5-7B $+20.40$ $-23.58$ $+44.20$
DeepSeek-R1-Distill-Qwen-7B $+18.40$ $-25.38$ $+27.80$
Qwen2.5-72B $+18.80$ $-22.38$ $+34.80$
BERT $+44.60$ $-22.78$ $+33.60$
RoBERTa $+43.00$ $-23.18$ $+35.80$
Phi4-14B $+19.20$ $-23.98$ $+34.60$
Mistral-Large $+48.63$ $-21.78$ $+38.20$
User Study $80.00$ $62.22$ $100.00$
Qwen2.5-14B Qwen2.5-14B $+21.24$ $-28.62$ $+17.80$
DeepSeek-R1-Distill-Qwen-14B $+20.64$ $-29.02$ $+18.40$
Qwen2.5-72B $+17.04$ $-28.02$ $+21.20$
BERT $+46.04$ $-23.02$ $+31.80$
RoBERTa $+43.84$ $-24.02$ $+38.60$
Phi4-14B $+14.84$ $-27.62$ $+13.80$
Mistral-Large $+27.04$ $-28.42$ $+22.00$
User Study $84.44$ $57.78$ $100.00$
Qwen2.5-32B Qwen2.5-32B $-10.33$ $-22.13$ $+30.00$
DeepSeek-R1-Distill-Qwen-32B $-0.53$ $-22.33$ $+25.60$
Qwen2.5-72B $-16.21$ $-21.73$ $+23.80$
BERT $+9.62$ $-19.53$ $+33.00$
RoBERTa $+27.62$ $-15.13$ $+38.20$
Phi4-14B $-3.98$ $-22.93$ $+15.80$
Mistral-Large $+10.22$ $-18.13$ $+22.60$
User Study $62.22$ $66.67$ $100.00$
Llama3-8B Llama3-8B $+15.36$ $-23.53$ $+27.00$
DeepSeek-R1-Distill-Llama-8B $+17.36$ $-27.13$ $+26.80$
Llama3-70B $+22.56$ $-25.53$ $+20.80$
BERT $+36.56$ $-15.33$ $+32.80$
RoBERTa $+37.56$ $-23.13$ $+35.80$
Phi4-14B $+15.56$ $-23.53$ $+26.80$
Mistral-Large $+23.56$ $-21.13$ $+32.40$
User Study $75.56$ $46.67$ $100.00$
Llama3-70B Llama3-70B $+1.84$ $-43.38$ $+16.40$
DeepSeek-R1-Distill-Llama-70B $+3.84$ $-41.70$ $+18.40$
Llama3-8B $-1.76$ $-40.58$ $+26.20$
BERT $+18.24$ $-36.98$ $+31.40$
RoBERTa $+21.24$ $-36.78$ $+37.40$
Phi4-14B $+0.84$ $-42.38$ $+22.80$
Mistral-Large $+11.64$ $-31.38$ $+29.80$
User Study $64.44$ $42.22$ $100.00$
(a) Counterfactual examples generated using FIZLE.
Model Judge Model ($LLM_{\mathcal{J}}$) AG News SST2 SNLI
Qwen2.5-7B Qwen2.5-7B $+44.58$ $-24.08$ $+31.72$
DeepSeek-R1-Distill-Qwen-7B $+37.78$ $-26.49$ $+41.09$
Qwen2.5-72B $+40.78$ $-27.75$ $+32.63$
BERT $+48.98$ $-23.28$ $+32.93$
RoBERTa $+45.78$ $-24.42$ $+29.00$
Phi4-14B $+41.38$ $-25.91$ $+31.42$
Mistral-Large $+24.78$ $-21.10$ $+49.85$
User Study $77.78$ $53.33$ $100.00$
Qwen2.5-14B Qwen2.5-14B $46.93$ $-22.44$ $+32.29$
DeepSeek-R1-Distill-Qwen-14B $+42.13$ $-24.16$ $+46.71$
Qwen2.5-72B $+41.93$ $-22.44$ $+36.05$
BERT $+53.53$ $-16.93$ $+30.09$
RoBERTa $+53.13$ $-19.00$ $+33.23$
Phi4-14B $+41.13$ $-22.44$ $+33.23$
Mistral-Large $+40.73$ $-12.46$ $+54.55$
User Study $73.33$ $66.67$ $100.00$
Qwen2.5-32B Qwen2.5-32B $+38.31$ $-35.36$ $+32.54$
DeepSeek-R1-Distill-Qwen-32B $+37.71$ $-39.49$ $+50.36$
Qwen2.5-72B $+35.91$ $-37.54$ $+34.68$
BERT $+47.91$ $-31.92$ $+30.64$
RoBERTa $+47.41$ $-32.26$ $+34.92$
Phi4-14B $+36.71$ $-36.39$ $+30.64$
Mistral-Large $+38.91$ $-24.92$ $+53.21$
User Study $71.11$ $51.11$ $100.00$
Llama3-8B Llama3-8B $+50.27$ $11.29$ $+34.86$
DeepSeek-R1-Distill-Llama-8B $+54.87$ $-1.10$ $+40.46$
Llama3-70B $+64.47$ $+4.29$ $+29.26$
BERT $+68.27$ $+8.77$ $+31.04$
RoBERTa $+68.07$ $+6.24$ $+29.77$
Phi4-14B $+56.07$ $+3.61$ $+28.50$
Mistral-Large $+58.47$ $+9.34$ $+62.34$
User Study $86.67$ $73.33$ $100.00$
Llama3-70B Llama3-70B $+39.13$ $-25.05$ $+36.52$
DeepSeek-R1-Distill-Llama-70B $+34.73$ $-27.11$ $+44.58$
Llama3-8B $+26.53$ $-12.09$ $+33.25$
BERT $+42.13$ $-22.53$ $+30.48$
RoBERTa $+39.73$ $-23.67$ $+37.53$
Phi4-14B $+33.53$ $-20.69$ $+37.78$
Mistral-Large $+37.53$ $-11.63$ $+47.60$
User Study $73.33$ $62.22$ $100.00$
(b) Counterfactual examples generated using FLARE.
Table 1: The LFR difference ($\Delta\%$) between the user study and the judge–generated model relationships (with values closer to 0 indicating better alignment). Counterfactual examples are generated by Qwen2.5-7B,14B,32B and Llama3-8B,70B, evaluated across judge models exhibiting same, distilled, same family, and independent w/ and w/o fine-tuning relationships on AG News, SST2 and SNLI. The user study to assess the LFR is conducted on 45 selected counterfactuals (§4.5). Orange-highlighted values indicate that the judge model with the given relationship aligns closely with human annotators, while purple-highlighted values indicate the opposite.

### v2

@@ Abstract
Counterfactual examples are widely employed to enhance the performance and robustness of large language models (LLMs) through counterfactual data augmentation (CDA). However, the selection of the judge model used to evaluate label flipping, the primary metric for assessing the validity of generated counterfactuals for CDA, yields inconsistent results. To decipher this, we define four types of relationships between the counterfactual generator and judge models: being the same model, belonging to the same model family, being independent models, and having an distillation relationship. Through extensive experiments involving two state-of-the-art LLM-based methods, three datasets, four generator models, and 15 judge models, complemented by a user study $(n=90)$, we demonstrate that judge models with an independent, non-fine-tuned relationship to the generator model provide the most reliable label flipping evaluations. [^1] Code and evaluation results are available at: https://github.com/qiaw99/failure Relationships between the generator and judge models, which are closely aligned with the user study for CDA, result in better model performance and robustness. Nevertheless, we find that the gap between the most effective judge models and the results obtained from the user study remains considerably large. This suggests that a fully automated pipeline for CDA may be inadequate and requires human intervention.
@@ 1 Introduction
Counterfactual examples are minimally altered versions of original inputs that flip the initial label Miller (2019); Ross et al. (2021); Madsen et al. (2022). They serve as a valuable approach for CDA aimed at improving model robustness and performance Liu et al. (2021); Dixit et al. (2022); Balashankar et al. (2023); Agrawal et al. (2025). We want to emphasize the subtle yet significant distinction between counterfactuals used for explaining model predictions and those used for CDA. In the former, the objective is to flip the model’s prediction, while the goal of CDA is to flips the ground truth label (Figure 6 in Appendix B). [^2] In this study, we mainly focus on the latter type of counterfactuals used for CDA. To evaluate the effectiveness and validity of the LLM-generated counterfactuals for CDA, the label flip rate (LFR) is a common metric of choice Ge et al. (2021). It is the percentage of valid counterfactuals where the ground truth labels are flipped out of the total number of instances. LFR of counterfactuals can be evaluated using either the same model that generates the counterfactual Bhattacharjee et al. (2024a, b); Wang et al. (2025a) or independent models Dixit et al. (2022); Balashankar et al. (2023). The optimal strategy for selecting models to evaluate the ground-truth validity of counterfactuals remains uncertain (Figure 1). This uncertainty, in turn, hampers efforts to enhance model robustness and performance through CDA, as noisy or erroneous labels may degrade model performance.
Figure 1: A counterfactual generated by Llama3-8B, with its label evaluated by judge models with different relationships, complemented by human evaluation. The revised words are highlighted in red.
@@
In this work, we first define four types of relationships between the counterfactual generator model and the judge model: being the same model, independent models with and without fine-tuning on the target dataset, distilled models, and models from the same family (Figure 1). Secondly, we conduct comprehensive experiments to predict labels for counterfactuals generated by two state-of-the-art approaches across three datasets and four generator models, with 15 judge models. Thirdly, we undertake a user study to assess the validity of the generated counterfactuals in acquiring ground-truth LFR. Our finding shows that a judge model with an independent, non-fine-tuned relationship to the generator captures label flipping most effectively. Relationships between the generator and judge models that are most aligned with the user study, lead to improved model performance and robustness. Additionally, there remains a considerable gap between the performance of the best judge models and the results observed in the user study.
2 Background and Related Work
CDA Pipeline
Initially, LLMs generate counterfactuals using established counterfactual generation approaches (Figure 4 in Appendix A). These generated counterfactuals should subsequently be validated—either by human annotators or judge models—as invalid counterfactuals bearing incorrect labels may degrade model performance Song et al. (2023). In practice, relying solely on human evaluation is both costly and inefficient; therefore, LLMs are commonly employed to validate the generated counterfactuals. Finally, valid counterfactuals are utilized as additional training data to enhance model performance and robustness Yang et al. (2021); Dixit et al. (2022).
@@ 2 Background and Related Work
The literature identifies two primary ways to verify label flipping in edited inputs: (1) employing an independent model, distinct from the one producing counterfactuals; (2) utilizing the same LLM that generates counterfactuals, guided by carefully constructed prompts. Prior to, and even during the widespread adoption of LLMs, encoder-only models, e.g., DeBERTa Dixit et al. (2022), RoBERTa Ross et al. (2021); Balashankar et al. (2023); Treviso et al. (2023) or BERT Kaushik et al. (2020); Fern and Pope (2021); Robeer et al. (2021), were predominantly used to verify label flipping. This preference stems from their advanced performance in text classification tasks. In more recent work, the same LLMs are increasingly employed both to generate counterfactuals Bhattacharjee et al. (2024a, b); Wang et al. (2025a); Dehghanighobadi et al. (2025) and to evaluate them for label flipping, since their classification accuracy rivals that of fine-tuned encoder-only models.
@@ 3.1 Label Flipping
Given that counterfactual examples used for CDA are defined as edited inputs that alter ground truth labels, LFR is positioned as the primary evaluation metric for assessing the effectiveness and validity of generated counterfactuals Kaushik et al. (2020); Dixit et al. (2022); Zhu et al. (2023); Chen et al. (2023). LFR is quantified as the percentage of instances in which labels are successfully flipped relative to the total number of counterfactuals, where $N$ stands for the total number of counterfactuals, $y_{k}$ represents the ground-truth label of the original input, $y_{k}^{{}^{\prime}}$ denotes the prediction of its corresponding counterfactual and $\mathds{1}$ is the indicator function:
@@ 3.2 Relationships
To determine the optimal model selection strategy for label flip evaluation, we identified four prevalent relationships $\mathcal{R}(\mathcal{G},\mathcal{J})$ between the generator model $LLM_{\mathcal{G}}$ and judge models $LLM_{\mathcal{J}}$ (Figure 1), which are used to assess label flipping:
  • Same model ($\mathcal{R}_{sm}$): the two models are same.
    $LLM_{\mathcal{G}}=LLM_{\mathcal{J}}$
@@
    $LLM_{\mathcal{G}},LLM_{\mathcal{J}}\in\mathcal{F(M)}$
  • Independent models ($\mathcal{R}_{im}$): the two models belong to different model families.
    $LLM_{\mathcal{G}}\in\mathcal{F}(\mathcal{M}_{1}),LLM_{\mathcal{J}}\in\mathcal{F}(\mathcal{M}_{2})$
    We further distinguish the independent models based on whether $LLM_{\mathcal{J}}$ is fine-tuned on the given dataset: independent models with ($\mathcal{R}_{imw}$) and without ($\mathcal{R}_{imwo}$) fine-tuning.
@@
    $LLM_{\mathcal{D}}$ $\notin\mathcal{F(M)}$
    $\text{Size}(LLM_{\mathcal{J}})$ $=\text{Size}(LLM_{\mathcal{G}})$
    $\text{Archit.}(LLM_{\mathcal{J}})$ $=\text{Archit.}(LLM_{\mathcal{G}})$
    $LLM_{\mathcal{J}}$ $=\text{Inherit}(LLM_{\mathcal{D}})$
    $\{LLM_{\mathcal{G}},LLM_{\mathcal{J}}\}$ $\cap LLM_{\mathcal{D}}=\varnothing$
@@ 4.1 Counterfactual Methods Selection
We select two state-of-the-art approaches based on LLMs that are shown to generate counterfactual examples efficiently and effectively: FIZLE Bhattacharjee et al. (2024a) and FLARE Bhattacharjee et al. (2024b). FIZLE first enables LLMs to identify key words within the input and then leverages these words to guide the generation of counterfactual examples. Meanwhile, FLARE generates counterfactuals by prompting LLMs in three steps: extracting latent features, identifying relevant words linked to those features, and modifying these words to produce counterfactual examples.
@@ 4.2 Datasets
We use three widely studied classification tasks for counterfactual generation in the literature [^3] Examples of the dataset and the label distribution are included in Appendix B.: news topic classification, sentiment analysis, and natural language inference.
@@ 4.3 Models
We select four LLMs varying in parameter size – Qwen2.5-{14B,32B} Qwen (2024), Llama3-{8B,70B} AI@Meta (2024) – to generate counterfactuals and serve as judge models $LLM_{\mathcal{J}}$ as $\mathcal{R}_{sm}$ relationship (§3.2). Additionally, we deploy fine-tuned BERT Devlin et al. (2019) and RoBERTa Liu et al. (2020) on the target datasets (§4.2), along with off-the-shelf Phi4-14B Abdin et al. (2024), Qwen2.5-72B Qwen (2024), Mistral-Large-Instruct Jiang et al. (2023), DeepSeek-R1-Distill-{Qwen,Llama} DeepSeek-AI (2025), and Gemini-1.5-pro Gemini (2024) as $LLM_{\mathcal{J}}$ [^4] Detailed information about the models employed is provided in Appendix C, and the downstream task performance of each model across three datasets and the classification prompts used are presented in Appendix D. (Table 3). We further ensemble label flipping results from all judge models via majority voting to yield final labels (ensemble in Figure 2). Moreover, since $LLM_{\mathcal{J}}$ are used to identify label flipping, we evaluate their downstream task performance in terms of classification accuracy across three datasets: $LLM_{\mathcal{J}}$ with $\mathcal{R}_{imw}$ relationship (BERT and RoBERTa) generally achieve the highest downstream performance (Appendix D).
4.4 User Study
We recruit 90 native English speakers and, for each of the three datasets, randomly sample 45 indices. For each subset, i.e., generator-dataset pair (Table 3 in Appendix G), the counterfactuals generated by the corresponding generator model $LLM_{\mathcal{G}}$ for the selected indices are evaluated by two human annotators. If no majority label emerges from the labels provided by human annotators, we determine one of the two labels chosen by the two annotators, ensuring the ground-truth label is agreed upon by two people. [^5] The annotation guidelines and annotator information are provided in Appendix E. We further conduct an automatic evaluation in Appendix G.3 on the selected 45 counterfactuals as a sanity check to validate their representativeness of the overall distribution. Each annotator is given 15 counterfactuals, along with the set of possible labels given by the dataset, and tasked with selecting the optimal label. We report an inter-annotator agreement of Cohen’s $\kappa=0.55$. [^6] Although we employ two state-of-the-art methods – FIZLE and FLARE – to generate counterfactuals, they are not consistently perfect (at times failing to fully shift the semantics from the original to the target label) and sometimes produce ambiguous cases (Figure 12). This has an effect on the IAA being moderate which would likely improve with the development and use of more advanced counterfactual-generation methods. Lastly, we calculate the ground-truth LFR as the proportion of valid counterfactuals relative to the total number of instances (Table 4 in Appendix G).
@@ 4.5.1 Counterfactual LFR Evaluation
LFR is evaluated based on the classification results of counterfactual examples (§3.1) generated using FIZLE and FLARE (§4.1), using the deployed judge models $LLM_{\mathcal{J}}$ described in Section 4.3.
4.5.2 Human Alignment Evaluation
To assess the alignment of $LLM_{\mathcal{J}}$ with human annotators, we employ three measures: (1) the average ranking ($rank\downarrow$, Figure 2); (2) the ratio of most-to-least alignment ($r_{m/l}\uparrow$); (3) Pearson correlation $\rho$ between human evaluation results and LFR results by judge models.
Average Ranking
To obtain the rankings, we first calculate LFR for human annotators and for each judge-generator model pair on each set of counterfactuals generated by given generator models, as reported in Table 4 of Appendix G. Next, we compute the difference ($\Delta$) between the human LFR and the LFR of each pair. A smaller difference indicates a better ranking (lower ranking value). This process results in a ranking for each judge-generator pair. Since these pairs correspond to specific relationships $\mathcal{R}(\mathcal{G},\mathcal{J})$, we average the rankings of all pairs sharing the same relationship to obtain the average ranking per relationship, as shown in Table 3 of Appendix G. Finally, we average these rankings across each generator model to produce the overall average rankings presented in Figure 2.
Most-to-Least-Alignment
Instead of measuring overall alignment, $r_{m/\ell}$ reports how many times each relationship $\mathcal{R}_{{\mathcal{G},\mathcal{J}}}$ most or least closely aligned with human annotators across the three datasets:
$r_{m/\ell}(\mathcal{R})=\frac{1}{|\mathcal{D}|}\sum_{d\in\mathcal{D}}\frac{\mathcal{N}_{\text{max}}(\mathcal{R},d)}{\mathcal{N}_{\text{min}}(\mathcal{R},d)}$
where $\mathcal{D}$ is the set of datasets and $\mathcal{N}_{\text{max}}$ and $\mathcal{N}_{\text{min}}$ denote the number of cases, in which a generator-judge model pair is the most closely and least aligned with human evaluation results, respectively. A higher $r_{m/l}$ reflects better alignment.